# Strange quark matter as dark matter: 40 years later, a reappraisal

### v1

@@ 5 Implications for cosmology
Interactions between dark matter particles and baryons generate a drag force that suppresses the growth of cosmic structures at small scales. This can be used to derive constraints on the dark matter-baryon scattering cross section. Dvorkin et al. (2014) consider a velocity-dependent cross section $\sigma_{\mathrm{scatt}}\propto v^{n}$, with $-4\leq n\leq 2$. For a velocity-independent cross section, they find $\sigma_{\mathrm{scatt}}/M<3.3\times 10^{-3}\,\mathrm{cm^{2}}\,\mathrm{g}^{-1}$ at 95% CL from a combination of CMB data from Planck and Ly-$\alpha$ data from the Sloan Digital Sky Survey. As in the case of dark matter self-interactions, these constraints are practically irrelevant for the clusters of SQM discussed in this paper.
@@ 6 Search for strangelets
  • Light Mass ($A\lesssim 10^{5}$) , quasi-relativistic particles. These particles can be directly detected in the flux of cosmic rays as they cross space-borne or balloon-borne detectors. In the 1990 balloon experiment HECRO-81 Saito et al. (1990) reported the observation of two events with $Z\simeq 14$ and an estimated mass $A\simeq 350$. In 1987, the ARIEL-6 satellite (see Fowler et al. (1987)) employing Cherenkov counters, presented an analysis of $Z\geq 34$ finding no SQM candidates in 427 days of data acquisition. The HEAO-3 apparatus (see Binns et al. (1989)), with an exposure of $8\times 10^{11}$ $\mathrm{cm^{2}\>sr\>s}$, did not find any strangelet candidate. Among the passive detectors (similar to the ground mica experiments) the SkyLab experiment (Shirk & Price (1978)), with 1.2 $\mathrm{m}^{2}$ Lexan track detectors, did not find any valid candidate in the superheavy ($Z>110$) nuclei range. On board the space station Mir, the experiment TREK (Price et al. (1992a)) explored the $Z>50$ region, also finding no strangelet candidate. In the case of a magnetic spectrometer, strangelets would appear as events having a low deflection due to their high mass/charge (A/Z) ratio, due to the presence of the negative charge of the strange quarks (in case of Z=0 they would be undetectable). BESS balloon-borne detector flights have yielded no candidates for $5\leq Z\leq 26$ for $A/Z>5$ Ichimura et al. (1993). Subsequently, data from 2006 to 2009 of the PAMELA space spectrometer were employed to conduct a direct search for SQM within cosmic rays (Adriani et al., 2015a). The spectrometer explored a rigidity range from $1-1000$ GV without finding traces of such particles, thereby setting an upper limit on the monochromatic strangelet flux in cosmic rays for particles having charge $1\leq Z\leq 8$ and baryonic mass $4\leq A\leq 1.2\times 10^{5}$. The AMS-01 Shuttle-Mir flight reported the observation of two events: $Z=8$, $A=20$ (3.93 GV) and $Z=4$, $A=50$ (5.13 GV), probably due to background, since the following, much longer, AMS-02 mission has found no events and has lowered the limit to the flux of $Z=2$, yielding the limits represented in Fig. 7.
@@
  • Underground detector, Intermediate mass $10^{-11}\,\mathrm{g}<M\leq 1\,\mathrm{g}$. Underground, the MACRO detector placed a limit of $1.1\times 10^{-14}$ cm-2 sr-1 s-1 and $5.5\times 10^{-15}$ cm-2 sr-1 s-1in the mass ranges $10^{-11}\,\mathrm{g}<m<0.1\,\mathrm{g}$ and $m>0.1\,\mathrm{g}$ respectively Ahlen et al. (1991),
@@
  • High Mass, Meteor-like particle $10^{-6}\,\mathrm{g}\lesssim M\lesssim 10^{4}\,\mathrm{g}$. Since strangelets can be rather massive, their presence can be inferred through the seismic activity they can potentially induce on the Moon, as suggested in Banerdt et al. (2006). According to Nakamura & Frohlich (2006), most of a rare class of lunar seismic events detected by the Apollo lunar seismic network occurred when the Moon faced a specific direction of the celestial sphere, suggesting that they may be caused by objects of extra-solar origin. Strangelets can also give a meteor-like signal as they burn in the atmosphere, with tracks that are either longer and/or faster than those of conventional meteors. Pi in the SKY, searching for fast meteors in the sky (Piotrowski et al., 2020), yielded a limit close to the dark matter constraint. The DIMS experiment is currently searching for these events from the ground, using two cameras for stereoscopic view of meteors and determination of their trajectory (Abe et al., 2021). Similarly, Mini-EUSO is searching for these events from the ISS (Adams et al., 2015; Bacholle et al., 2021) with a wider field. Another detector, SQM-ISS, aiming for direct detection of SQM, is being designed to be placed on the ISS in the next years. In this case, the idea is to use a local compact detector composed of a stack of scintillators (to detect ionization by passing particles) and a stack of metal plates (to read the vibration caused by the passage of particles) coupled to a time-of-flight system capable of measuring the speed of the particle. Strangelets or similar particles are expected to move through the detector unaffected by the ISS hull, at speeds consistent with the galactic orbital velocity of $\sim$ 220 km/s or lower.
@@
  • High-Mass Asteroid-size celestial bodies with $M\lesssim 10^{21}\,\mathrm{g}$, such as Polyhymnia and Ludmilla (Carry (2012)) both with mass densities of about $75$ g/cm3 have been proposed as being constituted of either ultra-heavy (Z>120) nuclei or SQM. Although more prosaically their anomalous density is due to uncertainties in the mass determination related to the difficulty of orbital perturbations, a systematic search for anomalous density asteroids could place a limit in this mass range.
@@ 7 Conclusions
In this article we have revisited the possibility that dark matter is made of clusters of up, down and strange quarks, as suggested by Witten (1984). We have shown that it is at least conceptually possible to formulate a scheme, leading to the formation of those objects in the Early Universe, without assuming the existence of a first-order phase transition associated with confinement. The recent experimental results, showing that nuclei can form at high temperatures in ultra-relativistic heavy ion collisions, support the claim that clusterization does not need to be associated with a first-order transition.
We have investigated a simple scheme, based on the formation of quark clusters at the hadronization temperature, which later partially evaporate, and we have shown that these objects can satisfy all the astrophysical and cosmological constraints on dark matter.
@@ Acknowledgements
This material is based upon work supported by the National Science Foundation under grants No. PHY-2208724, PHY-1654219 and PHY-2116686, and within the framework of the MUSES collaboration, under grant number No. OAC- 2103680. This material is also based upon work supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics, under Award Number DE-SC0022023 and by the National Aeronautics and Space Agency (NASA) under Award Number 80NSSC24K0767.

### v3

@@ 5 Implications for cosmology
Interactions between dark matter particles and baryons generate a drag force that suppresses the growth of cosmic structures at small scales. This can be used to derive constraints on the dark matter-baryon scattering cross section. Dvorkin et al. (2014) consider a velocity-dependent cross section $\sigma_{\mathrm{scatt}}\propto v^{n}$, with $-4\leq n\leq 2$. For a velocity-independent cross section, they find $\sigma_{\mathrm{scatt}}/M<3.3\times 10^{-3}\,\mathrm{cm^{2}}\,\mathrm{g}^{-1}$ at 95% C.L. from a combination of CMB data from Planck and Ly-$\alpha$ data from the Sloan Digital Sky Survey. As in the case of dark matter self-interactions, these constraints are practically irrelevant for the clusters of SQM discussed in this paper.
@@ 6 Search for strangelets
  • Light Mass ($A\lesssim 10^{5}$) , quasi-relativistic particles. These particles can be directly detected in the flux of cosmic rays as they cross space-borne or balloon-borne detectors. In the 1990 balloon experiment HECRO-81 (Saito et al., 1990) reported the observation of two events with $Z\simeq 14$ and an estimated mass $A\simeq 350$. In 1987, the ARIEL-6 satellite (see Fowler et al. (1987)) employing Cherenkov counters, presented an analysis of $Z\geq 34$ finding no SQM candidates in 427 days of data acquisition. The HEAO-3 apparatus (see Binns et al. (1989)), with an exposure of $8\times 10^{11}$ $\mathrm{cm^{2}\>sr\>s}$, did not find any strangelet candidate. Among the passive detectors (similar to the ground mica experiments) the SkyLab experiment (Shirk & Price, 1978), with 1.2 $\mathrm{m}^{2}$ Lexan track detectors, did not find any valid candidate in the superheavy ($Z>110$) nuclei range. On board the space station Mir, the experiment TREK (Price et al., 1992a) explored the $Z>50$ region, also finding no strangelet candidate. In the case of a magnetic spectrometer, strangelets would appear as events having a low deflection due to their high mass/charge (A/Z) ratio, due to the presence of the negative charge of the strange quarks (in case of Z=0 they would be undetectable). BESS balloon-borne detector flights have yielded no candidates for $5\leq Z\leq 26$ for $A/Z>5$ (Ichimura et al., 1993). Subsequently, data from 2006 to 2009 of the PAMELA space spectrometer were employed to conduct a direct search for SQM within cosmic rays (Adriani et al., 2015a). The spectrometer explored a rigidity range from $1-1000$ GV without finding traces of such particles, thereby setting an upper limit on the monochromatic strangelet flux in cosmic rays for particles having charge $1\leq Z\leq 8$ and baryonic mass $4\leq A\leq 1.2\times 10^{5}$. The AMS-01 Shuttle-Mir flight reported the observation of two events: $Z=8$, $A=20$ (3.93 GV) and $Z=4$, $A=50$ (5.13 GV), probably due to background, since the following, much longer, AMS-02 mission has found no events and has lowered the limit to the flux of $Z=2$, yielding the limits represented in Fig. 8.
@@
  • Underground detector, Intermediate mass $10^{-11}\,\mathrm{g}<M\leq 1\,\mathrm{g}$. Underground, the MACRO detector placed a limit of $1.1\times 10^{-14}$ cm-2 sr-1 s-1 and $5.5\times 10^{-15}$ cm-2 sr-1 s-1in the mass ranges $10^{-11}\,\mathrm{g}<m<0.1\,\mathrm{g}$ and $m>0.1\,\mathrm{g}$ respectively (Ahlen et al., 1991),
@@
  • High Mass, Meteor-like particle $10^{-6}\,\mathrm{g}\lesssim M\lesssim 10^{4}\,\mathrm{g}$. Since strangelets can be rather massive, their presence can be inferred through the seismic activity they can potentially induce on the Moon, as suggested in Banerdt et al. (2006). According to Nakamura & Frohlich (2006) most of a rare class of lunar seismic events detected by the Apollo lunar seismic network occurred when the Moon faced a specific direction of the celestial sphere, suggesting that they may be caused by objects of extra-solar origin. Strangelets can also give a meteor-like signal as they burn in the atmosphere, with tracks that are either longer and/or faster than those of conventional meteors. Pi in the SKY, searching for fast meteors in the sky (Piotrowski et al., 2020), yielded a limit close to the dark matter constraint. The DIMS experiment is currently searching for these events from the ground, using two cameras for stereoscopic view of meteors and determination of their trajectory (Abe et al., 2021). Similarly, Mini-EUSO is searching for these events from the ISS (Adams et al., 2015; Bacholle et al., 2021) with a wider field. Another detector, SQM-ISS, aiming for direct detection of SQM, is being designed to be placed on the ISS in the next years. In this case, the idea is to use a local compact detector composed of a stack of scintillators (to detect ionization by passing particles) and a stack of metal plates (to read the vibration caused by the passage of particles) coupled to a time-of-flight system capable of measuring the speed of the particle. Strangelets or similar particles are expected to move through the detector unaffected by the ISS hull, at speeds consistent with the galactic orbital velocity of $\sim$ 220 km/s or lower.
@@
  • High-Mass celestial bodies. Asteroid-size objects having a mass $M\lesssim 10^{21}\,\mathrm{g}$, such as Polyhymnia and Ludmilla (Carry, 2012) both with mass densities of about $75$ g/cm3 have been proposed as being constituted of either ultra-heavy (Z>120) nuclei or SQM. Although more prosaically their anomalous density is due to uncertainties in the mass determination related to the difficulty of orbital perturbations, a systematic search for anomalous density asteroids could place a limit in this mass range.
    A systematic search for objects in the mass range $10^{17}\,\mathrm{g}\lesssim M\lesssim 10^{23}\,\mathrm{g}$ could be based on their gravitational effects on the orbits of planets or satellites. For instance, it has been shown in Tran et al. (2024) that they can affect the orbit of Mars. Notice that the lower limit of $10^{17}\mathrm{g}$, indicated above, is due to the Hawking radiation as discussed in subsection 2.7. This limit applies to PBHs but not to strangelets which can have a smaller mass. Since direct observations of PBHs and strangelets are generally challenging, distinguishing between the two candidates can be difficult. However, detecting objects having a mass smaller than $10^{17}\,\mathrm{g}$ could offer a way to distinguish the two scenarios. It could be interesting to check if strangelets belonging to this mass range could be detected by an improved version of the Lunar Laser Ranging experiments (Dickey et al., 1994; Williams et al., 2004).
@@ 7 Conclusions
In this article we have revisited the possibility that dark matter is made of clusters of up, down and strange quarks, as suggested by Witten (1984). We have shown that it is at least conceptually possible to formulate a scheme, leading to the formation of those objects in the Early Universe, without assuming the existence of a first-order phase transition associated with confinement. In our analysis we have considered the possibility of producing massive strangelets through a first-order phase transition taking place in the high-temperature hadronic matter. A first possibility is based on the coalescence of sexaquarks (Shahrbaf et al., 2022). Another possibility is instead related to the existence of a chemical-mechanical instability leading to the production of high-density, finite-strangeness matter (Lavagno & Pigato, 2022). These theoretical possibilities need to be further investigated. We have then discussed how strangelets undergo a partial evaporation, and we have shown that these objects can satisfy all the astrophysical and cosmological constraints on dark matter. This result clearly depends on the value of a few phenomenological parameters, for instance the one related to the effect of Pauli blocking on the evaporation rate. It will be important to investigate if the value of these parameters can be justified by microphysics.
@@ Acknowledgements
This material is based upon work supported by the National Science Foundation under grants No. PHY-2208724, PHY-1654219 and PHY-2116686, and within the framework of the MUSES collaboration, under grant number No. OAC- 2103680. This material is also based upon work supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics, under Award Number DE-SC0022023 as well as by the National Aeronautics and Space Agency (NASA) under Award Number 80NSSC24K0767.
Data Availability
The data underlying this article will be shared on reasonable request to the corresponding author.